OpenStudy (anonymous):

x 2+5x+6 quadratic formula

OpenStudy (espex):

You could factor this to get your roots, unless you are expected to use the quadratic formula.

OpenStudy (espex):

Remember that your equation is in the form, \(ax^2+bx+c=0\) and your quadratic formula is \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Simply plug and play.